Real GDP is $8000, Autonomous consumption is $500, and planned investment is $200. MPC is .8. Given this income-expenditure equilibrium, why will firms tend to decrease output?
Answer #1

Find the aggregate expenditure using

AE = C + I  

= 500 + 0.8Y + 200

AE = 700 + 0.8Y

Equilibrium GDP is Y = 700 + 0.8Y

Y = 3500

This shows that equilibrium GDP and aggregate demand for goods and services is only 3500 but firms are producing a level of output equal to 8000. Because of excess supply and unintended inventory investment, firm will like to decrease output to reach equilibrium